# Accelerating Reionization Constraints: An ANN-Emulator Framework for the SCRIPT Semi-numerical Model

### v1

@@ 1 Introduction
The Epoch of Reionization (EoR) marks the last major phase transition in the universe, during which the intergalactic medium (IGM) was transformed from a cold, neutral state into a hot, ionized one [barkanaBeginningFirstSources2001]. Despite substantial theoretical and observational progress, the detailed timeline and physical drivers of reionization remain uncertain. A new generation of high-redshift observations, from facilities such as the Precision Array for Probing the Epoch of Reionization (PAPER) [parsonsPRECISIONARRAYPROBING2010], Low-Frequency Array (LOFAR) [haarlemLOFARLOwFrequencyARray2013], Murchison Widefield Array (MWA) [tingayMurchisonWidefieldArray2013], Square Kilometre Array (SKA) [koopmansCosmicDawnEpoch2015], Hydrogen Epoch of Reionization Array (HERA) [deboerHydrogenEpochReionization2017], and the James Webb Space Telescope (JWST) [gardnerJamesWebbSpace2023], is already delivering, or will soon deliver, unprecedented constraints on the underlying astrophysical processes. Fully exploiting these data, however, requires efficient parameter inference based on numerical or semi-numerical simulations capable of modeling the coupled evolution of ionization, thermal history, and galaxy populations.
@@
In recent years, a variety of machine learning (ML) techniques have been proposed to accelerate parameter inference in cosmological studies. These methods have been applied to infer parameters directly from the 21 cm power spectrum [e.g., shimabukuroAnalysing21Cm2017, doussotImprovedSupervisedLearning2019, choudhuryExtracting21cmPower2022, shimabukuroEstimationIiBubble2022, choudhuryInferringIGMParameters2025], global signal [e.g., choudhuryExtracting21Cm2020, choudhuryUsingArtificialNeural2021], as well as from images [e.g., gilletDeepLearning21cm2019, kwonDeepLearningStudy21cm2020, neutschInferringAstrophysicsDark2022], and lightcones [e.g., zhaoSimulationbasedInferenceReionization2022, prelogovicMachineLearningAstrophysics2022]. In particular, ML-based emulators have emerged as powerful surrogates for computationally expensive simulations. Trained on a representative set of simulations (the training dataset), these emulators can efficiently interpolate simulation outputs across high-dimensional parameter spaces with remarkable accuracy, offering results at a fraction of the computational cost of traditional methods. Among the various ML approaches used to construct emulators, artificial neural networks (ANNs) have been particularly popular [e.g., schmitEmulationReionizationSimulations2018, jenningsEvaluatingMachineLearning2019, cohenEmulatingGlobal21cm2020, mondalTightConstraintsExcess2020, bevinsGlobalemuNovelRobust2021, bye21cmVAEVeryAccurate2022, tiwariImprovingConstraintsReionization2022, breitman21cmemuEmulator21cmfast2023, sikderEmulationCosmicDawn2024, lazareHERABoundXray2024, meriotLORELIDatabase212024, choudhuryInferringIGMParameters2025, mahidaANNBNNInferring2025, maityEmulatorbasedForecastingAstrophysics2025, gonzalez-hernandezUsingNeuralEmulators2025], while Gaussian processes (GPs) have also seen widespread use [e.g., kernEmulatingSimulationsCosmic2017, maityFastMethodReionization2023, choudhuryGPRbasedEmulatorSeminumerical2024, gharaConstrainingIntergalacticMedium2020]. More recently, advanced deep learning architectures have been explored, including convolutional autoencoders and vision transformer-based models for emulating reionization maps [e.g., chardinDeepLearningModel2019, postureCosmoUiTVisionTransformerUNet2025], as well as generative adversarial networks (GANs) for lightcone image generation [e.g., diaoMultifidelityEmulatorLargescale2025].
@@
However, a central challenge in emulator construction lies in generating an effective training dataset. In practice these training datasets are generated using outputs of simulations evaluated at parameter points sampled from the model’s parameter space. Common sampling strategies include simple approaches, such as uniform random and grid sampling [e.g., chardinDeepLearningModel2019, cohenEmulatingGlobal21cm2020, sikderEmulationCosmicDawn2024, meriotLORELIDatabase212024, choudhuryInferringIGMParameters2025, facchinettiNeuralNetworkEmulation2025, mahidaANNBNNInferring2025, gharaConstrainingIntergalacticMedium2020, maityEmulatorbasedForecastingAstrophysics2025, postureCosmoUiTVisionTransformerUNet2025]. Uniform random sampling is easy to implement and ensures no bias towards any subregion, but it can lead to uneven coverage: samples can cluster by chance or leave gaps, especially in high dimensions. This strategy is effective if one can afford a very large sample size and the emulated quantity varies fairly smoothly with the parameters. However, it becomes inefficient in high-dimensional parameter spaces, as many samples are required to adequately cover the space, increasing the chances of large unsampled regions and redundant sampling along certain directions. By contrast, a grid sampling approach places points at fixed intervals along each parameter axis avoiding the problem of uneven coverage. But, this approach also suffers from a severe drawback: because each parameter is held fixed on a lattice, many samples end up sharing the same value in one or more parameters. This wastes computation time and means slowly-varying directions are over-sampled. Moreover, the total number of grid points required to cover the space grows exponentially with the dimensionality, leading to poor scalability to high-dimensional parameter spaces.
@@
A widely used alternative strategy is Latin hypercube (LH) sampling [e.g., schmitEmulationReionizationSimulations2018, jenningsEvaluatingMachineLearning2019, tiwariImprovingConstraintsReionization2022, lazareHERABoundXray2024, cooperSimulationBasedInference2025, gonzalez-hernandezUsingNeuralEmulators2025, diaoMultifidelityEmulatorLargescale2025]. In LH sampling, each parameter’s range is divided into $N$ equal-probability intervals, and one sample is drawn in each interval for each parameter. The $N$ samples are arranged so that no two share the same value in any parameter. This ensures that each parameter is evenly covered and avoids duplicate marginal values. Crucially, this sampling strategy overcomes the scalability limitations of uniform random and grid sampling methods. While this strategy performs well when parameters are reasonably constrained and the priors are narrow, it becomes inefficient when parameters are poorly constrained and the priors are broad. In such cases, the emulated quantity may vary rapidly across the parameter space, requiring a large number of samples to adequately capture these variations. However, many of these samples will correspond to low-likelihood regions when used for inference, leading to significant computational waste and slower emulator development despite the increased sampling effort.
@@
A possible solution to the above problem was demonstrated in Kern et al. [kernEmulatingSimulationsCosmic2017], where they first employed coarse LH samples spanning a broad parameter range to build an approximate emulator and performed a preliminary MCMC exploration to locate the high-likelihood region of the parameter space. They then refined the emulator by adding denser, spherical training sets around the estimated maximum a posteriori point. A related idea appears in Breitman et al. [breitman21cmemuEmulator21cmfast2023], where they trained the emulator using samples drawn from a previous inference run, ensuring dense coverage of the high-likelihood region in parameter space. Although both strategies mitigate the inefficiency of sampling over broad priors by concentrating computational effort within regions of high-likelihood, they still have their drawbacks. In the first approach, a preliminary MCMC exploration using an emulator trained on coarse LH samples drawn from a broad prior can lead to spurious posterior distributions that have little to no overlap with the true posterior obtained from a full MCMC run. Consequently, the inferred high-likelihood region may be systematically misplaced, greatly reducing the reliability of subsequent refinements. In the second approach, while utilizing samples from previous inference runs ensures dense coverage of the high-likelihood region, it relies on the availability of existing posterior samples that accurately represent the model of interest. Such data may not exist for new or modified models, or for simulations with expanded parameter spaces, thereby limiting the general applicability of this method.
@@
In this work, we introduce an ANN-based likelihood emulator for the semi-numerical code SCRIPT [choudhuryPhotonNumberConservation2018, maityProbingThermalHistory2022]. This framework is designed to overcome the challenges associated with existing targeted sampling methods for training dataset generation by exploiting a key property of SCRIPT: its numerical convergence at coarse resolutions for large-scale properties relevant to observational constraints [maityFastMethodReionization2023]. This key property allows us to build a framework that directly overcomes the challenges of existing targeted sampling methods. First, to solve the potential unreliability of preliminary MCMC explorations built on coarse, non-convergent samples [kernEmulatingSimulationsCosmic2017], we leverage the convergence of SCRIPT to run a full, reliable MCMC using inexpensive coarse-resolution simulations. This ensures our training samples are drawn directly from the true high-likelihood region, as developed and implemented in previous works [maityFastMethodReionization2023, choudhuryGPRbasedEmulatorSeminumerical2024]. Second, to remain general-purpose and not rely on pre-existing inference runs [breitman21cmemuEmulator21cmfast2023], our method generates this training set from scratch.
@@
The remainder of this paper is organized as follows. In section 2, we describe the reionization model, outline the observational constraints, and present the results of a conventional MCMC analysis that serve as a benchmark. In section 3, we first provide a brief overview of ANNs, then describe generation of the training dataset and the training procedure, and finally compare the parameter constraints obtained from conventional MCMC with those from ANN-based emulators. We summarize our key findings in section 4. The cosmological parameters used in this work are $\Omega_{m}=0.308,\,\Omega_{\Lambda}=1-\Omega_{m},\,\Omega_{b}=0.0482,\,h=0.678,\,\sigma_{8}=0.829,\text{ and }n_{s}=0.961$ [adePlanck2015Results2016].
@@ 2 Conventional MCMC Analysis
In this section, we describe the reionization model, observational constraints, and results of conventional MCMC-based parameter inference analyses. These will serve as the benchmark for comparing our emulator-based parameter constraints in the next section. This section draws closely on previous works [maityConstrainingReionizationThermal2022, maityFastMethodReionization2023], and is included for completeness.
@@ 2.1 The reionization model
The model used in this work is based on the semi-numerical photon conserving framework SCRIPT [choudhuryPhotonNumberConservation2018], which provides the ionization state of the universe in a cosmologically representative simulation volume. The main feature of this framework is its ability to explicitly conserve the ionizing photons and hence produce numerically convergent power spectra of ionization fluctuations with respect to the resolution of the ionization maps. In particular, we use a five-parameter SCRIPT model that is extended to incorporate inhomogeneous recombinations and to compute the thermal history, thereby enabling the self-consistent inclusion of radiative feedback effects [maityProbingThermalHistory2022, maityConstrainingReionizationThermal2022, maityFastMethodReionization2023]. For a given redshift, we supply our model with the density field and the distribution of collapsed halos capable of producing ionizing radiation. The model then produces the corresponding ionization field at that redshift. To generate the density fields, we use the second-order Lagrangian perturbation theory (2LPT) approximation rather than a full N-body simulation, as our focus is solely on capturing the large-scale features of the intergalactic medium (IGM). In particular, we use GADGET-2 [springelCosmologicalSimulationCode2005a] plugins provided by the 2LPT density field generator MUSIC [hahnMultiscaleInitialConditions2011] to generate the density fields. For computing the distribution of collapsed halos, we use a sub-grid prescription based on the conditional ellipsoidal mass function [shethExcursionSetModel2002]. In this work, we use a simulation box size of $256~h^{-1}$ cMpc, which is large enough for the observables that we are interested in computing. We use simulation boxes with redshifts from $z=5$ to $20$ with an interval of $\Delta z=0.1$, which enables us to compute the reionization history.
@@
Our model requires the ionization efficiency $\zeta$, which is used to compute the ionization maps by estimating the number of available ionizing photons per hydrogen atom. To include inhomogeneous recombinations, we tune the ionization criteria to compensate for excess neutral atoms, and incorporate small-scale fluctuations through a globally averaged clumping factor $C_{\mathrm{HII}}$, fixing its value $C_{\mathrm{HII}}=3$ motivated from earlier simulation studies [daloisioHydrodynamicResponseIntergalactic2020]. The model also solves for the thermal history of each grid cell in the box and accounts for effects of spatially inhomogeneous reionization on the thermal evolution by assuming that a region’s temperature increments by a value $T_{\mathrm{re}}$, known as the reionization temperature, as it is ionized for the first time. In addition, our model also incorporates radiative feedback by suppressing production of ionizing photons in halos where the gas is heated up. Several methods of incorporating this feedback have been introduced in [maityProbingThermalHistory2022]. In this work, we use the “step feedback” model.
@@
Our model has five free parameters as described below:
  1. (i)
@@
  3. (iii)
    Finally, the model needs the escape fraction $f_{\mathrm{esc}}$, which is the fraction of photons that escape from halos of mass $M_{h}$, for calculating the UV luminosity functions (UVLFs). Since we only use UVLF data from $z=6$ and $7$ as observational constraints, we ignore its $z$ dependence and parameterize the escape fraction as
@@ 2.2 Observational constraints
The observational constraints used in this work are similar to those used in our previous works [maityConstrainingReionizationThermal2022, maityFastMethodReionization2023]. These are summarized below:
  1. (i)
    We use the latest Planck measurement of the CMB optical depth $\tau_{e}=0.054\pm 0.007$ [aghanimPlanck2018Results2020].
  2. (ii)
    We employ model-independent lower bounds on the ionized fraction derived from the dark-pixel analysis of quasar spectra by McGreer et al. [mcgreerModelindependentEvidenceFavour2015]. These measurements provide lower limits on the ionization fraction at $z=5.6$ and $z=5.9$. More recent analyses [jinNearlyModelindependentConstraints2023, zhuProbingUltralateReionization2023, daviesUpdatedDarkPixel2025] have updated these constraints using improved datasets, in some cases tightening the limits and in others relaxing them. For methodological continuity and comparability with our earlier work, however, we adopt the values reported in ref. [mcgreerModelindependentEvidenceFavour2015].
  3. (iii)
@@
  4. (iv)
    We utilize data on galaxy UVLFs at $z=6$ and $7$ obtained from optical studies [bouwensUVLUMINOSITYFUNCTIONS2015, bouwens6LuminosityFunction2017]. We do not incorporate the latest JWST UVLF measurements in this work. The main goal of this work is to develop and validate an ANN-emulator framework for parameter inference and to benchmark its performance against that of a conventional MCMC. Accurately modelling the evolving galaxy UVLFs over the full redshift range $6\lesssim z\lesssim 15$ requires a more sophisticated galaxy formation model, which in turn introduces a significantly larger number of free parameters [chakrabortyModellingStarformationActivity2024, choudhuryCapturingSmallScaleReionization2025]. This added complexity renders an MCMC-based parameter inference using high-resolution simulations computationally infeasible, preventing a fair comparison with our ANN-based results. In future, we aim to extend our framework to a fourteen-parameter model [choudhuryCapturingSmallScaleReionization2025], incorporating the latest UVLFs.
  5. (v)
@@
  6. (vi)
    Finally, we exclude models that produce reionization histories exhibiting “double reionization”, in which the ionized fraction undergoes a significant decrease after an initial rise before reaching unity again. Such scenarios are incompatible with our temperature-evolution model, which assumes photoionization equilibrium; this assumption breaks down if regions recombine more rapidly than they are reionized.

### v2

@@ 1 Introduction
The Epoch of Reionization (EoR) marks the last major phase transition in the universe, during which the intergalactic medium (IGM) was transformed from a cold, neutral state into a hot, ionized one [3]. Despite substantial theoretical and observational progress, the detailed timeline and physical drivers of reionization remain uncertain. A new generation of high-redshift observations, from facilities such as the Low-Frequency Array (LOFAR) [82], Murchison Widefield Array (MWA) [78], Square Kilometre Array (SKA) [43], Hydrogen Epoch of Reionization Array (HERA) [25], Atacama Large Millimeter/submillimeter Array (ALMA) [83], and the James Webb Space Telescope (JWST) [30], is already delivering, or will soon deliver, unprecedented constraints on the underlying astrophysical processes. Fully exploiting these data, however, requires efficient parameter inference based on numerical or semi-numerical simulations capable of modeling the coupled evolution of the ionization state and thermal history of the IGM, together with galaxy populations.
@@
In recent years, a variety of machine learning (ML) techniques have been proposed to accelerate parameter inference in cosmological studies. These methods have been applied to infer parameters directly from the 21 cm power spectrum [e.g., 72, 27, 16, 71, 17], global signal [e.g., 15, 14], as well as from images [e.g., 34, 44, 61], and lightcones [e.g., 84, 66]. In particular, ML-based emulators have emerged as powerful surrogates for computationally expensive simulations. Trained on a representative set of simulations (the training dataset), these emulators can efficiently interpolate simulation outputs across high-dimensional parameter spaces with remarkable accuracy, offering results at a fraction of the computational cost of traditional methods. Among the various ML approaches used to construct emulators, artificial neural networks (ANNs) have been particularly popular [e.g., 69, 40, 21, 60, 4, 9, 79, 8, 73, 45, 58, 17, 52, 56, 36], while Gaussian processes (GPs) have also seen widespread use [e.g., 42, 55, 18, 33]. More recently, advanced deep learning architectures have been explored, including convolutional autoencoders and vision transformer-based models for emulating reionization maps [e.g., 11, 65], as well as generative adversarial networks (GANs) for lightcone image generation [e.g., 26].
@@
However, a central challenge in emulator construction lies in generating an effective training dataset. In practice these training datasets are generated using outputs of simulations evaluated at parameter points sampled from the model’s parameter space. Common sampling strategies include simple approaches, such as uniform random and grid sampling [e.g., 11, 21, 73, 58, 17, 28, 52, 33, 56, 65]. Uniform random sampling is easy to implement and ensures no bias towards any subregion, but it can lead to uneven coverage: samples can cluster by chance or leave gaps, especially in high dimensions. This strategy is effective if one can afford a very large sample size and the emulated quantity varies fairly smoothly with the parameters. However, it becomes inefficient in high-dimensional parameter spaces, as many samples are required to adequately cover the space, increasing the chances of large unsampled regions and redundant sampling along certain directions. By contrast, a grid sampling approach places points at fixed intervals along each parameter axis avoiding the problem of uneven coverage. But, this approach also suffers from a severe drawback: because each parameter is held fixed on a lattice, many samples end up sharing the same value in one or more parameters. This wastes computation time and means slowly-varying directions are over-sampled. Moreover, the total number of grid points required to cover the space grows exponentially with the dimensionality, leading to poor scalability to high-dimensional parameter spaces.
@@
A widely used alternative strategy is Latin hypercube (LH) sampling [e.g., 69, 40, 79, 45, 22, 36, 26]. In LH sampling, each parameter’s range is divided into $N$ equal-probability intervals, and one sample is drawn in each interval for each parameter. The $N$ samples are arranged so that no two share the same value in any parameter. This ensures that each parameter is evenly covered and avoids duplicate marginal values. Crucially, this sampling strategy overcomes the scalability limitations of uniform random and grid sampling methods. While this strategy performs well when parameters are reasonably constrained and the priors are narrow, it becomes inefficient when parameters are poorly constrained and the priors are broad. In such cases, the emulated quantity may vary rapidly across the parameter space, requiring a large number of samples to adequately capture these variations. However, many of these samples will correspond to low-likelihood regions when used for inference, leading to significant computational waste and slower emulator development despite the increased sampling effort.
@@
A possible solution to the above problem was demonstrated in Kern et al. [42], where they first employed coarse LH samples spanning a broad parameter range to build an approximate emulator and performed a preliminary MCMC exploration to locate the high-likelihood region of the parameter space. They then refined the emulator by adding denser, spherical training sets around the estimated maximum a posteriori point. A related idea appears in Breitman et al. [8], where they trained the emulator using samples drawn from a previous inference run, ensuring dense coverage of the high-likelihood region in parameter space. Although both strategies mitigate the inefficiency of sampling over broad priors by concentrating computational effort within regions of high-likelihood, they still have their drawbacks. In the first approach, a preliminary MCMC exploration using an emulator trained on coarse LH samples drawn from a broad prior can lead to spurious posterior distributions that have little to no overlap with the true posterior obtained from a full MCMC run. Consequently, the inferred high-likelihood region may be systematically misplaced, greatly reducing the reliability of subsequent refinements. In the second approach, while utilizing samples from previous inference runs ensures dense coverage of the high-likelihood region, it relies on the availability of existing posterior samples that accurately represent the model of interest. Such data may not exist for new or modified models, or for simulations with expanded parameter spaces, thereby limiting the general applicability of this method.
@@
In this work, we introduce an ANN-based likelihood emulator for the semi-numerical code SCRIPT [20, 54]. This framework is designed to overcome the challenges associated with existing targeted sampling methods for training dataset generation by exploiting a key property of SCRIPT: its numerical convergence at coarse resolutions for large-scale properties relevant to observational constraints [55]. This key property allows us to build a framework that directly overcomes the challenges of existing targeted sampling methods. First, to solve the potential unreliability of preliminary MCMC explorations built on coarse, non-convergent samples [42], we leverage the convergence of SCRIPT to run a full, reliable MCMC using inexpensive coarse-resolution simulations. This ensures our training samples are drawn directly from the true high-likelihood region, as developed and implemented in previous works [55, 18]. Second, to remain general-purpose and not rely on pre-existing inference runs [8], our method generates this training set from scratch.
@@
The remainder of this paper is organized as follows. In section 2, we describe the reionization model, outline the observational constraints, and present the results of a conventional MCMC analysis that serve as a benchmark. In section 3, we first provide a brief overview of ANNs, then describe generation of the training dataset and the training procedure, and finally compare the parameter constraints obtained from conventional MCMC with those from ANN-based emulators. We summarize our key findings in section 4. The cosmological parameters used in this work are $\Omega_{m}=0.308,\,\Omega_{\Lambda}=1-\Omega_{m},\,\Omega_{b}=0.0482,\,h=0.678,\,\sigma_{8}=0.829,\text{ and }n_{s}=0.961$ [63].
@@ 2 Conventional MCMC analysis
In this section, we describe the reionization model, observational constraints, and results of conventional MCMC-based parameter inference analyses. These will serve as the benchmark for comparing our emulator-based parameter constraints in the next section. This section draws closely on previous works [53, 55], and is included for completeness.
@@ 2.1 The reionization model
The model used in this work is based on the semi-numerical photon conserving framework SCRIPT [20], which provides the ionization state of the universe in a cosmologically representative simulation volume. The main feature of this framework is its ability to explicitly conserve the ionizing photons and hence produce numerically convergent power spectra of ionization fluctuations with respect to the resolution of the ionization maps. In particular, we use a five-parameter SCRIPT model that is extended to incorporate inhomogeneous recombinations and to compute the thermal history, thereby enabling the self-consistent inclusion of radiative feedback effects [54, 53, 55]. For a given redshift, we supply our model with the density field and the distribution of collapsed halos capable of producing ionizing radiation. The model then produces the corresponding ionization field at that redshift. To generate the density fields, we use the second-order Lagrangian perturbation theory (2LPT) approximation rather than a full N-body simulation, as our focus is solely on capturing the large-scale features of the intergalactic medium (IGM). In particular, we use GADGET-2 [75] plugins provided by the 2LPT density field generator MUSIC [37] to generate the density fields. For computing the distribution of collapsed halos, we use a sub-grid prescription based on the conditional ellipsoidal mass function [70]. In this work, we use a simulation box size of $256~h^{-1}$ cMpc, which is large enough for the observables that we are interested in computing. We use simulation boxes with redshifts from $z=5$ to $20$ with an interval of $\Delta z=0.1$, which enables us to compute the reionization history.
@@
Our model requires the ionization efficiency $\zeta$, which is used to compute the ionization maps by estimating the number of available ionizing photons per hydrogen atom. To include inhomogeneous recombinations, we tune the ionization criteria to compensate for excess neutral atoms, and incorporate small-scale fluctuations through a globally averaged clumping factor $C_{\mathrm{HII}}$, fixing its value $C_{\mathrm{HII}}=3$ motivated from earlier simulation studies [23]. The model also solves for the thermal history of each grid cell in the box and accounts for effects of spatially inhomogeneous reionization on the thermal evolution by assuming that a region’s temperature increments by a value $T_{\mathrm{re}}$, known as the reionization temperature, as it is ionized for the first time. In addition, our model also incorporates radiative feedback by suppressing production of ionizing photons in halos where the gas is heated up. Several methods of incorporating this feedback have been introduced in [54]. In this work, we use the “step feedback” model.
In addition to the ionization history, we compute the galaxy UV luminosity function (UVLF) by relating the ionizing properties of the sources of reionization to their non-ionizing UV emission. The ionization efficiency $\zeta$ is a multiplicative combination of the star formation efficiency $f_{\star}$ and the escape fraction of ionizing photons $f_{\mathrm{esc}}$. While the UVLF is sensitive only to $f_{\star}$, computing it from $\zeta$ requires independent knowledge of $f_{\mathrm{esc}}$ (or equivalently $f_{\star}$) in order to disentangle these two quantities. An additional ingredient required to compute the UVLF is the relation between the ionizing and non-ionizing UV properties of the stellar population. We specify this using the slope of the galaxy spectrum above the Lyman limit, $\alpha_{s}$, and the ratio of luminosities at the Lyman continuum (corresponding to rest wavelength 912 Å) and at 1500 Å, $\mathcal{R}_{912/\mathrm{UV}}\equiv L_{912}/L_{\mathrm{UV}}$. As in previous works [54, 53], we adopt $\alpha_{s}=2$ and $\mathcal{R}_{912/\mathrm{UV}}=0.2$, consistent with galaxies having a Salpeter IMF and metallicity $Z\approx 0.1\,Z_{\odot}=0.002$, as estimated using Starburst99 [46]. We note that both these stellar population parameters are degenerate with the escape fraction. We also note that the UV continuum is modulated by dust attenuation [e.g., 12, 31, 51]. However, because this effect is largely degenerate with the star formation efficiency $f_{\star}$, we do not attempt to model dust attenuation separately in our analysis.
@@
Within this framework, our model is described by five free parameters as described below:
  1. 1.
@@
  3. 3.
    Finally, as described above, the model needs the escape fraction $f_{\mathrm{esc}}$, which is the fraction of photons that escape from halos of mass $M_{h}$, for calculating the UVLFs. Since we only use UVLF data from $z=6$ and $7$ as observational constraints, we ignore its $z$ dependence and parameterize the escape fraction as
@@
    Because $\zeta\propto f_{\star}f_{\mathrm{esc}}$, assuming a halo mass independent ionization efficiency $\zeta$ while introducing a halo mass dependence in the escape fraction of the form $f_{\mathrm{esc}}\propto M_{h}^{\beta}$ implies an effective inverse halo mass dependence for the star formation efficiency, $f_{\star}\propto M_{h}^{-\beta}$. We emphasize that this should be interpreted as an effective scaling arising from the adopted parametrization, rather than as an assumption that the star formation efficiency is intrinsically independent of halo mass. Indeed, numerous studies have shown that feedback-regulated star formation leads to a strong halo mass dependence of $f_{\star}$, particularly at $z\sim 6$77, which plays a key role in shaping the UVLF [e.g., 76, 77, 74]. Within this effective framework, the resulting scaling of $f_{\star}$ is broadly consistent with parameter estimates obtained in earlier studies [e.g., 62, 67]. However, we note that this assumption is adopted here as a simplifying choice. In the future, as additional observational data become available, we plan to constrain an extended version of our model [19], in which both $f_{\mathrm{esc}}$ and $f_{\star}$ are allowed to have independent halo mass dependencies.
@@ 2.2 Observational constraints
The observational constraints used in this work are similar to those used in our previous works [53, 55]. These are summarized below:
  1. 1.
    We use the latest Planck measurement of the CMB optical depth $\tau_{e}=0.054\pm 0.007$ [64].
  2. 2.
    We employ model-independent lower bounds on the ionized fraction derived from the dark-pixel analysis of quasar spectra by McGreer et al. [57]. These measurements provide lower limits on the ionization fraction at $z=5.6$ and $z=5.9$. More recent analyses [41, 86, 24] have updated these constraints using improved datasets, in some cases tightening the limits and in others relaxing them. For methodological continuity and comparability with our earlier work, however, we adopt the values reported in ref. [57].
  3. 3.
@@
  4. 4.
    We utilize data on galaxy UVLFs at $z=6$ and $7$ obtained from optical studies [6, 7]. We do not incorporate the latest JWST UVLF measurements in this work. The main goal of this work is to develop and validate an ANN-emulator framework for parameter inference and to benchmark its performance against that of a conventional MCMC. Accurately modeling the evolving galaxy UVLFs over the full redshift range $6\lesssim z\lesssim 15$ requires a more sophisticated galaxy formation model, which in turn introduces a significantly larger number of free parameters [10, 19]. This added complexity renders an MCMC-based parameter inference using high-resolution simulations computationally infeasible, preventing a fair comparison with our ANN-based results. In future, we aim to extend our framework to a fourteen-parameter model [19], incorporating the latest UVLFs.
  5. 5.
    We include estimates of the thermal state of the low-density IGM. These are parameterized using the temperature-density relation, which for low-density gas can be well approximated by a power-law relation
    $T=T_{0}\Delta^{\gamma-1},$ (2.3)
@@
  6. 6.
    Finally, we exclude models that produce reionization histories exhibiting “double reionization”, in which the ionized fraction undergoes a significant decrease after an initial rise before reaching unity again. Such scenarios are incompatible with our temperature-evolution model, which assumes photoionization equilibrium; this assumption breaks down if regions recombine more rapidly than they are reionized. In our parametrization, these histories arise for extreme negative values of $\alpha$, which cause the ionization efficiency $\zeta(z)$ to be very high at high redshifts and to decline sharply toward lower redshifts. In combination with a moderate value of $\zeta_{0}$, this can lead to an early phase of efficient ionization, followed by a decrease in the ionization fraction due to recombinations as $\zeta(z)$ decreases.